# PointAD+: Learning Hierarchical Representations for Zero-shot 3D Anomaly Detection

### v2

@@ Abstract
Zero-shot (ZS) 3D anomaly detection is a critical yet underexplored field that addresses scenarios where target 3D training samples are unavailable. This presents a challenge, as it requires the model to accurately detect 3D anomalies in previously unseen categories. In this paper, we aim to transfer CLIP’s robust 2D generalization capabilities to identify 3D anomalies across unseen objects of highly diverse class semantics. To this end, we propose a unified framework to comprehensively detect and segment 3D anomalies by leveraging both point- and pixel-level information. We first design PointAD, which leverages point-pixel correspondence to represent 3D anomalies through their associated rendering pixel representations. This approach is referred to as implicit 3D representation, as it focuses solely on rendering pixel anomalies but neglects the inherent spatial relationships within point clouds. Then, we propose PointAD+ to further broaden the interpretation of 3D anomalies by introducing explicit 3D representation, emphasizing spatial abnormality to uncover abnormal spatial relationships. Hence, we propose G-aggregation to involve geometry information to enable the aggregated point representations spatially aware. To simultaneously capture rendering and spatial abnormality, PointAD+ proposes hierarchical representation learning, incorporating implicit and explicit anomaly semantics into hierarchical text prompts: rendering prompts for the rendering layer and geometry prompts for the geometry layer. A cross-hierarchy contrastive alignment is further introduced to promote the interaction between the rendering and geometry layers, facilitating mutual anomaly learning. Finally, PointAD+ integrates anomaly semantics from both layers to capture the generalized anomaly semantics. During the test, PointAD+ can integrate RGB information in a plug-and-play manner and further improve its detection performance. Extensive experiments demonstrate the superiority of PointAD+ in ZS 3D anomaly detection across unseen objects with highly diverse class semantics, achieving a holistic understanding of abnormality.
@@
Anomaly detection, zero-shot anomaly detection, 3D anomaly detection, vision-language model.
@@ 1 Introduction
This paper proposes a unified framework to transfer the 2D knowledge of CLIP to 3D understanding for ZS 3D anomaly detection. We first present PointAD (conference version), which understands 3D anomalies by capturing rendering abnormality: (1) deriving 2D representations of point clouds via CLIP by rendering them from multiple views, (2) obtaining 3D prediction by aggregating the 2D predictions according to the correspondence between points and pixels. Building on this, we propose PointAD+ to further enhance the interpretation of 3D anomalies by introducing spatial abnormality. The joint learning of rendering and spatial abnormality enables PointAD+ to achieve a holistic understanding of 3D anomalies.
As illustrated in Figure LABEL:fig2:PointAD, PointAD models 3D abnormality by leveraging point-pixel relationships, capturing 3D abnormality from both 2D and 3D representations. Since pixel anomalies are 2D projections from 3D point anomalies, we refer to them as implicit 3D abnormality. These pixel anomalies contribute to point anomaly modeling from two perspectives: (1) pixels in different views reflect partial information of the same point; (2) pixels corresponding to the same point should exhibit similar semantics. After this, we introduce hybrid representation learning to optimize the rendering prompts from both 2D and 3D for implicit point representation. Specifically, to fully capture view-specific representations for 3D anomaly understanding, we treat each view as a separate 2D anomaly detection task. These 2D tasks across multiple views jointly facilitate the point anomaly learning from a multi-task learning (MTL) perspective. On the other hand, since pixel representations from the same point can be regarded as different instances, we aggregate their prediction results via multiple instance learning (MIL) to represent 3D anomalies.
@@
As illustrated in Figure LABEL:fig2:PointAD+, PointAD+ instead characterizes 3D abnormality in two aspects: implicit 3D abnormality on the rendering layer and explicit 3D abnormality on the geometry layer. In addition to implicit 3D abnormality, PointAD+ further introduces explicit 3D abnormality to uncover abnormal spatial relationships, offering a comprehensive interpretation of 3D anomalies. Holding these two aspects of 3D anomalies, we propose hierarchical representation learning to incorporate both implicit and explicit 3D anomaly semantics into the learnable text prompts in two folds. On the rendering layer, we use hybrid representation learning, as utilized in PointAD, for implicit point learning. On the geometry layer, we propose explicit point representation learning that reveals abnormal spatial relations by incorporating geometric information. To achieve this, we propose G-aggregation, which explicitly aggregates 2D representations from different views and incorporates geometric information by representing point features via their neighboring features in 3D space. This allows the aggregated 3D representations to become spatially aware and effectively model abnormal spatial relationships. After capturing 3D abnormalities through both explicit and implicit point learning, we consider the anomaly semantics on the rendering and geometry layers to stem from the same root cause. As a result, these layers can complement and learn from one another. Hence, we establish a cross-hierarchy contrastive alignment that brings the anomaly semantics from both layers together while pushing normal and anomalous semantics apart. This facilitates mutual anomaly learning across both hierarchies. During the test, we fuse the implicit and explicit 3D anomaly semantics for comprehensive anomaly modeling.
Benefiting from the alignment between point and pixel anomaly semantics, PointAD and PointAD+ can seamlessly incorporate additional RGB information and perform ZS multimodal 3D (M3D) detection without the need for extra modules or retraining. Extensive experiments across datasets with 32 class semantics demonstrate the superiority of PointAD and PointAD+ on ZS 3D and M3D anomaly detection and segmentation, in both one-class and cross-dataset settings. PointAD+ shows the performance improvement compared to PointAD. We also provide a comprehensive analysis of key modules’ effectiveness and hyperparameters’ robustness, including rendering quality and lighting conditions. The main contributions of this paper are summarized as follows:
  • To the best of our knowledge, we take the first attempt to simultaneously investigate the challenging yet valuable ZS 3D and M3D anomaly detection. We propose to leverage the strong generalization of CLIP to detect and segment 3D anomalies over diverse unseen objects.
@@
  • Comprehensive experiments show that PointAD+ and PointAD outperform the baselines in ZS 3D and M3D by a large margin. We hope that our models will serve as a springboard for future research on ZS 3D and M3D anomaly detection.
@@ 2 Related Work
3D anomaly detection has recently gained attention in the community of anomaly detection. Unlike 2D anomaly detection, which leverages RGB information, 3D anomaly detection focuses on identifying unusual spatial relationships in 3D objects. While 2D anomaly detection has been extensively explored, research on 3D anomaly detection remains relatively limited. This disparity is largely due to the challenges associated with acquiring high-quality 3D data, as collecting 3D datasets is more costly and labor-intensive compared to 2D images. MVTec 3D-AD [20] is the first public dataset, which provides the point cloud anomalies and the corresponding 2D-RGB information. Subsequently, Eyecandies [39] and Real3D-AD [40] provide the 3D dataset for a wide exploration [41]. These datasets bridge the connection between 3D and 2D anomaly detection. 3D-ST [20] uses a teacher net to extract dense local geometry descriptors and design a student net to match such descriptors. AST [21] introduces asymmetric teacher and student net to further improve 3D anomaly detection. Instead of only using point clouds, BTF [19], M3DM [23, 42], CPFM [43], and SDM [25] integrate point features and RGB pixel features to detect 3D anomalies. While these approaches exhibit commendable performance by storing object-specific normal point and pixel features within the unsupervised learning framework, such paradigms simultaneously limit their generalization capacity to point clouds from unseen objects, which is crucial to detecting anomalies when the target object is unavailable. To fill this gap, we introduce PointAD+ and PointAD to identify unseen anomalies across diverse objects. They extend CLIP to the realm of ZS 3D and show robust generalization in capturing generic normality and abnormality within points. Furthermore, PointAD+ and PointAD serve as a unified framework, allowing seamless integration of 3D and RGB modalities without additional training.
@@ 2.3 Prompt Learning
With the increase in model parameters, training a model from scratch requires substantial computational resources and time. This high training cost hinders the rapid development of deep learning technologies. Therefore, rather than training or fine-tuning the entire network, prompt learning focuses on optimizing a minimal set of parameters to adapt the network for downstream tasks. CoOp, a seminal work in prompt learning [34, 51], introduces global context optimization to update learnable text prompts for few-shot recognition. DenseCLIP [35] extends this approach to dense classifications. More recently, AnomalyCLIP [16] introduces object-agnostic prompt learning to capture generic normality and abnormality in images for 2D anomaly detection. Other works, such as [52] and [53], have built upon this by incorporating additional vision information. In contrast, PointAD+ innovatively transforms the flat prompts used in these methods into hierarchical prompts to capture both rendering and spatial abnormality for ZS 3D anomaly detection. Leveraging this hierarchical structure, we introduce hierarchy representation learning to achieve an in-depth understanding of ZS 3D and M3D anomaly detection.
@@ 3 PointAD+
3.1 A Review of CLIP
@@ 3.2 Overview of PointAD and PointAD+
ZS 3D anomaly detection requires the model to identify anomalies in previously unseen 3D objects with diverse semantic features. Our goal is to transfer the strong 2D generalization capabilities of CLIP from images to 3D point cloud recognition. In this paper, we propose a unified framework for detecting and segmenting 3D anomalies in a ZS manner. As illustrated in Figure 3, PointAD+ understands point clouds from both pixel and point perspectives. To enable CLIP to comprehend 3D point clouds, we first render point clouds from multiple views and utilize CLIP’s visual encoder to extract pixel representations from these 2D renderings. We then aggregate the pixel representations corresponding to the same point to derive point-level representations. Based on this, we propose hierarchical representation learning to optimize the learnable hierarchical text prompts, allowing the model to incorporate generic anomaly semantics from both spatial abnormality (geometry layer) and rendering abnormality (rendering layer). On the rendering layer, we introduce hybrid representation learning to model rendering abnormality, taking into account regularization from both individual views and interactions between views (PointAD). On the geometry layer, we propose G-aggregation to explicitly obtain point representations by incorporating the geometry information within point clouds. Built on these two layers, we further facilitate the overall understanding of PointAD+ via knowledge sharing across rendering and geometry layers. To this end, we introduce cross-hierarchy alignment to promote mutual anomaly learning across different layers. Finally, we integrate the anomaly semantics from both hierarchies to generate the final outputs. Additionally, with explicit alignment between 3D and 2D representations, PointAD+ and PointAD can directly incorporate 2D RGB information during testing to enable ZS M3D detection.
@@
Figure 3: Framework of PointAD+. PointAD+ interprets 3D abnormality through implicit and explicit 3D abnormality. For implicit 3D abnormality, CLIP’s vision encoder extracts 2D global and local representations from the renderings, and then the resulting 2D representations are projected into 3D point representations to capture point anomaly semantics (rendering layer). For explicit 3D abnormality, we propose G-aggregation to incorporate geometric information into the aggregated point representations, balancing the specificity of individual points with their spatial continuity (geometry layer). Holding these layers, cross-hierarchy alignment is further introduced to facilitate mutual learning across layers. Finally, hierarchical representation learning jointly optimizes the text embeddings to align explicit and implicit point representations with learnable rendering and geometric prompts, capturing generic anomaly patterns comprehensively.
@@ 3.3 Multi-View Rendering
3.4 Pixel representation and point representation
Our framework aims to learn generic anomaly semantics from both pixel and point representations, enabling a comprehensive understanding of 3D anomalies. For a point cloud $x^{3d}_{i}$, we first obtain the 2D renderings $\mathcal{X}_{i}=\{x_{i}^{(k)}\}_{k=1}^{K}$. Then, these renderings are encoded via the vision encoder of CLIP to obtain global pixel representations $\mathcal{F}_{i}\hskip-1.99997pt=\hskip-1.99997pt\{f_{i}^{(k)}\}_{k=1}^{K}$, and local pixel representations $\mathcal{F}_{i}^{m}\hskip-1.99997pt=\hskip-1.99997pt\{f_{i}^{m(k)}\}_{k=1}^{K}$. As for point cloud representations, we consider two types of point representations, i.e., implicit point representation and explicit point representation. As for implicit point representation, the consistency constraint requires points and the corresponding pixels to maintain the same semantics. Therefore, The global implicit point representation $p$ and local implicit point representations ${p}^{m}$ are expected to include their corresponding pixel representations in each view. Formally, let us define the global implicit point representation of the $i$-thth point cloud as $p_{i}=\{p_{i}^{(k)}|p_{i}^{(k)}=f_{i}^{(k)}\}_{k=1}^{K}$. Also, we denote the local implicit point representation $p^{m}_{i}=\{p^{m(k)}_{i}|p^{m(k)}_{i}=\{p^{m(k)}_{i,j}\}_{j=1}^{n}\}_{k=1}^{K}$, where $p^{m(k)}_{i,j}=f_{i,u,v}^{m(k)}$ with $(u,v)=R^{\prime(k)}(a_{i,j},b_{i,j},c_{i,j})\}$. ($a_{i,j}$, $b_{i,j}$, $c_{i,j}$) represents the 3D coordinate of the $j$-thth point of $i$-thth point cloud. $R^{\prime(k)}$ is the transformation between the point and pixel space, derived as $R^{\prime(k)}=\frac{h}{H}R^{(k)}$.

### v3

# PointAD+: Learning Hierarchical Representations for ZS 3D Anomaly Detection

@@ Abstract
Zero-shot (ZS) 3D anomaly detection is a critical yet underexplored field that addresses scenarios where target 3D training samples are unavailable. This presents a challenge, as it requires the model to accurately detect 3D anomalies in previously unseen categories. In this paper, we aim to transfer CLIP’s robust 2D generalization capabilities to identify 3D anomalies across unseen objects of highly diverse class semantics. We first design PointAD, which leverages point-pixel correspondence to represent 3D anomalies through their associated rendering pixel representations. This representation is referred to as implicit point representation, as it focuses solely on rendering pixel anomalies but neglects the inherent spatial relationships within point clouds. However, spatial awareness modeling is important for accurate 3D abnormality recognition. Therefore, we propose PointAD+ to further broaden the interpretation of 3D anomalies by introducing explicit point representation that is aggregated from implicit point representation, modeling abnormal point relative position. We further introduce G-aggregation to involve geometry information to enable the explicit point representation to be spatially aware. To simultaneously capture rendering and geometry abnormality, PointAD+ proposes hierarchical representation learning, incorporating implicit and explicit anomaly semantics into hierarchical text prompts: rendering prompts for the rendering layer and geometry prompts for the geometry layer. A cross-hierarchy contrastive alignment is introduced to promote the interaction between the rendering and geometry layers, facilitating mutual anomaly learning. Finally, PointAD+ integrates anomaly semantics from both layers to capture the generalized anomaly semantics. During the test, PointAD+ can integrate RGB information in a plug-and-play manner and further improve its detection performance. Extensive experiments demonstrate the superiority of PointAD+ in ZS 3D anomaly detection across unseen objects with highly diverse class semantics, achieving a holistic understanding of abnormality.
@@
Anomaly detection, ZS anomaly detection, 3D anomaly detection, vision-language model.
@@ 1 Introduction
This paper proposes a unified framework to transfer the 2D knowledge of CLIP to 3D understanding for ZS 3D anomaly detection. We first present PointAD (conference version), which understands 3D anomalies by capturing rendering abnormality: (1) deriving 2D representations of point clouds via CLIP by rendering them from multiple views, (2) obtaining 3D prediction by aggregating the 2D predictions according to the correspondence between points and pixels. Specifically, as illustrated in Figure LABEL:fig2:PointAD, PointAD models 3D abnormality by leveraging point-pixel relationships. Since pixel anomalies are 2D projections from 3D point anomalies, we refer to them as implicit 3D abnormality. These pixel anomalies contribute to point anomaly modeling from two perspectives: (1) pixels in different views reflect partial information of the same point; (2) pixels corresponding to the same point should exhibit similar semantics. After this, we introduce hybrid representation learning to optimize the rendering prompts from both 2D and 3D for implicit point representation. Specifically, to fully capture view-specific representations for 3D anomaly understanding, we treat each view as a separate 2D anomaly detection task. These 2D tasks across multiple views jointly facilitate the point anomaly learning from a multi-task learning (MTL) perspective. On the other hand, since pixel representations from the same point can be regarded as different instances, we aggregate their prediction results via multiple instance learning (MIL) to represent 3D anomalies.
PointAD relies on view-wise 2D features to represent points indirectly. However, CLIP itself cannot capture the geometric information embedded in point clouds. Such spatial awareness is crucial for modeling accurate point representations, particularly for reliable abnormality recognition. PointAD+ instead characterizes 3D abnormality in two aspects: implicit 3D abnormality on the rendering layer and explicit 3D abnormality on the geometry layer. Holding these two aspects of 3D anomalies, we propose PointAD+ that learns hierarchical representations: (1) align implicit point representations on the rendering layer, which interpret each point through aggregated pixel-level logits from multi-view renderings; and (2) align explicit point representations on the geometry layer, aggregated from implicit point representations and explicitly encoding spatial geometry.
@@
In Figure LABEL:fig2:PointAD+, On the rendering layer, we use hybrid representation learning, as utilized in PointAD, for implicit point learning. On the geometry layer, we propose explicit point representation learning that reveals abnormal spatial relations by incorporating geometric information. To achieve this, we explicitly aggregate implicit point representations and propose G-aggregation to incorporate geometric information by representing point features via their neighboring features in 3D space. This allows the aggregated 3D representations to become spatially aware and effectively model abnormal spatial relationships. In addition, we consider the anomaly semantics on the rendering and geometry layers to stem from the same root cause. As a result, these layers can complement and learn from one another. Hence, we establish a cross-hierarchy contrastive alignment that brings the anomaly semantics from both layers together while pushing normal and anomalous semantics apart. This facilitates mutual anomaly learning across both hierarchies. During the test, we fuse the implicit and explicit 3D anomaly semantics for comprehensive anomaly modeling.
Benefiting from the alignment between point and pixel anomaly semantics, PointAD and PointAD+ can seamlessly incorporate additional RGB information and perform ZS multimodal 3D (multimodal 3D) detection without the need for extra modules or retraining. Extensive experiments across datasets with 32 class semantics demonstrate the superiority of PointAD and PointAD+ on ZS 3D and multimodal 3D anomaly detection and segmentation, in both one-class and cross-dataset settings. PointAD+ shows the performance improvement compared to PointAD. We also provide a comprehensive analysis of key modules’ effectiveness and hyperparameters’ robustness, including rendering quality and lighting conditions. The main contributions of this paper are summarized as follows:
  • To the best of our knowledge, we take the first attempt to simultaneously investigate the challenging yet valuable ZS 3D and multimodal 3D anomaly detection. We propose to leverage the strong generalization of CLIP to detect and segment 3D anomalies over diverse unseen objects.
@@
  • Comprehensive experiments show that PointAD+ and PointAD outperform the baselines in ZS 3D and multimodal 3D by a large margin. We hope that our models will serve as a springboard for future research on ZS 3D and multimodal 3D anomaly detection.
Discussion: Compared to the conference version, the main extended contributions of the journal conference are as follows: Field contribution: PointAD+ is the first to categorize point learning in ZS 3D anomaly detection into implicit point learning and explicit point learning. It highlights that CLIP-based frameworks fail to capture the geometric information of point clouds, which is crucial for accurate 3D anomaly detection; Technical contribution: PointAD+ proposes an effective approach to explicitly model point-wise representations from 2D features and incorporate geometric information into point representations. Then, PointAD+ introduces hierarchical representation learning to simultaneously align implicit and explicit point representations. PointAD+ achieves consistent improvements across all datasets in the paper; Research contribution: This work opens a new avenue for future research by encouraging the integration of stronger point-learning methods, such as PointNet-style or DGCNN-style architectures, to more effectively embed spatial relations into 2D representations for structural 3D anomalies.
@@ 2 Related Work
3D anomaly detection has recently gained attention in the community of anomaly detection. Unlike 2D anomaly detection, which leverages RGB information, 3D anomaly detection focuses on identifying unusual spatial relationships in 3D objects. While 2D anomaly detection has been extensively explored, research on 3D anomaly detection remains relatively limited. This disparity is largely due to the challenges associated with acquiring high-quality 3D data, as collecting 3D datasets is more costly and labor-intensive compared to 2D images. MVTec 3D-AD [20] is the first public dataset, which provides the point cloud anomalies and the corresponding 2D-RGB information. Subsequently, Eyecandies [39] and Real3D-AD [40] provide the 3D dataset for a wide exploration [41]. These datasets bridge the connection between 3D and 2D anomaly detection. 3D-ST [20] uses a teacher net to extract dense local geometry descriptors and design a student net to match such descriptors. AST [21] introduces asymmetric teacher and student net to further improve 3D anomaly detection. Instead of only using point clouds, BTF [19], multimodal 3DM [23, wang2025m3dm], CPFM [43], and SDM [25] integrate point features and RGB pixel features to detect 3D anomalies. While these approaches exhibit commendable performance by storing object-specific normal point and pixel features within the unsupervised learning framework, such paradigms simultaneously limit their generalization capacity to point clouds from unseen objects, which is crucial to detecting anomalies when the target object is unavailable. To fill this gap, we introduce PointAD+ and PointAD to identify unseen anomalies across diverse objects. They extend CLIP to the realm of ZS 3D and show robust generalization in capturing generic normality and abnormality within points. Furthermore, PointAD+ and PointAD serve as a unified framework, allowing seamless integration of 3D and RGB without additional training.
@@ 2.3 Prompt Learning
With the increase in model parameters, training a model from scratch requires substantial computational resources and time. This high training cost hinders the rapid development of deep learning technologies. Therefore, rather than training or fine-tuning the entire network, prompt learning focuses on optimizing a minimal set of parameters to adapt the network for downstream tasks. CoOp, a seminal work in prompt learning [34, 51], introduces global context optimization to update learnable text prompts for few-shot recognition. DenseCLIP [35] extends this approach to dense classifications. More recently, AnomalyCLIP [16] introduces object-agnostic prompt learning to capture generic normality and abnormality in images for 2D anomaly detection. Other works, such as [52] and [53], have built upon this by incorporating additional vision information. In contrast, PointAD+ innovatively transforms the flat prompts used in these methods into hierarchical prompts to capture both rendering and spatial abnormality for ZS 3D anomaly detection. Leveraging this hierarchical structure, we introduce hierarchy representation learning to achieve an in-depth understanding of ZS 3D and multimodal 3D anomaly detection.
@@ 3 PointAD+
(b) Visualization of explicit and implicit point score maps. Top row: point segmentation using explicit point learning; Middle row: point segmentation using implicit point learning; Bottom row: ground truths.
3.1 A Review of CLIP and AnomalyCLIP
@@
Replacing fixed text prompts with learnable text prompts is an effective method to adapt CLIP for downstream tasks. AnomalyCLIP [16] introduces object-agnostic text prompts to adapt CLIP for ZS 2D anomaly detection. The success of AnomalyCLIP demonstrates the promise of adapting CLIP for ZSAD by utilizing a lightweight, learnable text prompt. However, ZS 3D anomaly detection remains underexplored, primarily because there is no 3D foundation model with the strong generalization capabilities of CLIP. In this paper, we explore how to transfer CLIP’s strong generalization ability to 3D anomaly detection in a ZS manner.
@@ 3.2 Overview of PointAD and PointAD+
ZS AD requires models to identify anomalies in previously unseen 3D objects with diverse class semantics. Our goal is to transfer the strong 2D generalization ability of CLIP from images to 3D point cloud recognition. As illustrated in Figure 3, PointAD+ first interprets point clouds through (1) implicit point representations, which capture 3D abnormality via pixel features from 2D renderings (Section 3.5.1), and (2) explicit point representations, which aggregate implicit point representations for point-level representation and encode the geometry information via G-aggregation. (Section 3.5.2)
We then propose hierarchical representation learning (Section 3.6), which aligns these representations with learnable hierarchical text prompts, enabling the model to incorporate generic anomaly semantics from rendering-based abnormality (rendering/bottom layer) and geometry-based abnormality (geometry /top layer). On the rendering layer (Section 3.6.1), we introduce hybrid representation learning to align rendering prompts with both pixel representations from 2D renderings and implicit point representations in the implicit point space. On the geometry layer (Section 3.6.2), we align geometric prompts with explicit point representations, which are derived from implicit representations and further are encoded in geometry information, in explicit point space.
Built on these two layers, we further enhance the overall understanding of PointAD+ through cross-hierarchy alignment, which promotes mutual anomaly learning across rendering and geometry layers (Section 3.6.3). Finally, we integrate anomaly semantics from both hierarchies to generate the final outputs. Moreover, thanks to explicit alignment between 3D and 2D representations, both PointAD+ and PointAD can directly incorporate RGB information during testing, enabling ZS multi-modal 3D anomaly detection.
@@
Figure 3: Framework of PointAD+. PointAD+ interprets 3D abnormality through implicit and explicit 3D abnormality. For implicit 3D abnormality, CLIP’s vision encoder extracts 2D global and local representations from the renderings, and then the resulting 2D representations are projected into implicit point representation to capture point anomaly semantics (rendering layer). For explicit 3D abnormality, we propose G-aggregation to obtain the explicit point representation by aggregating the implicit point representation and then incorporating geometric information (geometry layer). Holding these layers, cross-hierarchy alignment is further introduced to facilitate mutual learning across layers. Finally, hierarchical representation learning jointly optimizes the text embeddings to align explicit and implicit point representations with learnable rendering and geometric prompts, capturing generic anomaly patterns comprehensively.
3.3 Motivation of PointAD+
PointAD employs implicit point representations to capture 3D anomaly semantics by aggregating pixel-level abnormalities into point-level ones. However, when anomalies are not well reflected in renderings, this approach can lead to missed anomalies and false alarms (Figure LABEL:fig:_ei1). In contrast, explicit point segmentation encodes geometric information to localize anomaly regions with higher precision, as it is more sensitive to spatial variations. Yet, explicit point learning is not universally superior: as shown in Figure LABEL:fig:_ei2, it struggles with edge points, where abrupt spatial changes often trigger false alarms. Implicit point representations, by relying on 2D features, are less sensitive to such edge variations and avoid this problem. To exploit the strengths of both, PointAD+ introduces hierarchical representation learning that integrates implicit point learning on the rendering layer and explicit point learning on the geometry layer, enabling a more comprehensive understanding of abnormalities.
@@ 3.4 Multi-view Rendering
3.5 Pixel and Point representation
PointAD+ interprets the 3D abnormality in implicit point space and explicit point space. In implicit point space, PointAD+ models pixel representation extracted by CLIP and implicit point representation that implicitly models point abnormality through point-pixel relationships. In explicit point space, PointAD+ explicitly models the point-level representation by aggregating implicit point representation and using G-aggregation to incorporate geometric information.
3.5.1 Implicit point space
Rendering abnormality in implicit point space reflects on pixel representation and implicit point representation.
Pixel representation: For a point cloud $x^{3d}_{i}$, we first obtain the 2D renderings $\mathcal{X}_{i}=\{x_{i}^{(k)}\}_{k=1}^{K}$. Then, these renderings are encoded via the vision encoder of CLIP to obtain pixel representation: global pixel representations $\mathcal{F}_{i}\hskip-1.99997pt=\hskip-1.99997pt\{f_{i}^{(k)}\}_{k=1}^{K}$, and local pixel representations $\mathcal{F}_{i}^{m}\hskip-1.99997pt=\hskip-1.99997pt\{f_{i}^{m(k)}\}_{k=1}^{K}$.
Implicit point representation: The consistency constraint requires points and the corresponding pixels to maintain the same semantics. Therefore, the global implicit point representation $p$ and local implicit point representations ${p}^{m}$ are expected to include their corresponding pixel representations in each view. Formally, let us define the global implicit point representation in the $i$-thth point cloud as $p_{i}=\{p_{i}^{(k)}|p_{i}^{(k)}=f_{i}^{(k)}\}_{k=1}^{K}$. Also, we denote the local implicit point representation:
$p^{m}_{i}=\{p^{m(k)}_{i}|p^{m(k)}_{i}=\{p^{m(k)}_{i,1},p^{m(k)}_{i,2},\cdots,p^{m(k)}_{i,n}\}\}_{k=1}^{K},$ (2)
where $p^{m(k)}_{i,j}=f_{i,u,v}^{m(k)}$ with $(u,v)=R^{\prime(k)}(a_{i,j},b_{i,j},c_{i,j})\}$. ($a_{i,j}$, $b_{i,j}$, $c_{i,j}$) represents the 3D coordinate of the $j$-thth point of $i$-thth point cloud. $R^{\prime(k)}$ is the transformation between the point and pixel space, derived as $R^{\prime(k)}=\frac{h}{H}R^{(k)}$.